OpenStudy (anonymous):

given sin theta = 2/3 and cos theta = sqrt 5/3, find csc theta

OpenStudy (anonymous):

\[\csc(\theta)=\frac{1}{\sin(\theta)}\] in other words, take the reciprocal of sine to find cosecant